A 400 mm x 400 mm column supports a DL = 500 KN, LL = 1000 KN, MDL = 600 KN-m, and MLL = 600 KN-m. It is required that the column must be placed at the center of the square footing. Assume that the weight of the footing and the soil above is to be 15% of the total service load. The soil bearing capacity is 209 KPa, fc’ = 27.6 MPa, and fy = 414 MPa. Design completely the square footing. Provide the footing plan and elevation. Use ACI Method
Answer #1

Given column size = 400mm x 400mm ~ 1611 X 1611 DL=500KN 7 113k LL = 1000 KN ~ 225k MDL=600kn-m= 442 kaft MLL = 600 KN-m = 44We know that Bearing pressure for strength aesign ( Total loa Алеа -14 (498 121 = 4.13 ksf ☆ calculation of depth required an117 calculation of depth required for one Way shear - (2.83+2) = 4.83 We know that Vu=1(2-82X4-1385 = 128.56 ksi bo = (1/12or 35 400 opsi -0.00316 60oOopsi As = 6.0033)x13 132 inch xzuinch = 10.45 in 2 over the entire 132 inch wiath use 11 bars of